Calculate the change in entropy when 1 mole nitrogen gas expands isothermally and reversibly from an initial volume of 1 litre to a final volume of 10 litre at 27 Degrees C. Thank you!


  1. Calculate the change in entropy when  1 mole nitrogen gas expands isothermally and reversibly from an initial volume of 1 litre to a final volume of 10 litre at 27 Degrees C.


Thank you!

Grade:11

2 Answers

Chetan Mandayam Nayakar
13 years ago

dS=dQ/T=nRTln(V2/V1)/T=8.314*ln10 J/K

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Change in entropy=2.303 RT logV2/v1

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi